# An ALMA spectroscopic survey of the Planck high-redshift object PLCK G073.457.5 confirms two protoclusters

### v1

@@
Planck’s High-Frequency Instrument observed the whole sky between 350 m and 3 mm, discovering thousands of unresolved peaks in the cosmic infrared background. The nature of these peaks are still poorly understood – while some are strong gravitational lenses, the majority are spatial overdensities of star-forming galaxies, but with almost no redshift constraints. PLCK G073.457.5 (G073) is one of these Planck-selected peaks. G073 was previously observed by ALMA, with the results suggesting the presence of two structures (one around redshift 1.5 and one around redshift 2) aligned along the line of sight, but the results lacked robust spectroscopic confirmation. Characterizing the full redshift distribution of the galaxies within G073 is needed in order to better understand this representative example of these Planck-selected objects, and connect them to the emergence of galaxy clusters. We use ALMA Band 3 spectral scans to search for CO(3–2), CO(4–3), and Ci(1–0) line emission, targeting eight red Herschel-SPIRE sources in the field, as well as four bright SCUBA-2 sources. We find 15 emission lines in 13 galaxies, and using existing photometry information we secure the spectroscopic redshifts of all 13 galaxies. 11 of these galaxies are SPIRE-selected and lie in two structures at and (both with a standard deviation in redshift of 0.02), while the two SCUBA-2-selected galaxies are at . Using multiwavelength photometry we constrain stellar masses and star-formation rates, and using the CO and Ci emission lines we constrain gas masses. Our protocluster galaxies exhibit typical depletion timescales (/SFR) compared to field galaxies at the same redshifts, but enhanced gas-to-stellar mass ratios, potentially driven by emission line selection effects. We find that the two structures confirmed in our survey are reproduced in cosmological simulations of star-forming halos at high redshift; the simulated halos have a 60–70% probability of collapsing into galaxy clusters, implying that the two structures in G073 are genuinely protoclusters.
@@
The Universe on the largest scales forms a ‘cosmic web’, with galaxy clusters at the intersections of the filaments (Bond et al., 1996). These are the largest gravitationally-bound objects in the Universe, and we believe that they play an important role on the evolution of the galaxies embedded within them because in the local Universe, cluster galaxies are more likely to be massive, red ellipticals with old stellar populations (e.g., Ellis et al., 1997; Andreon, 2003; Muzzin et al., 2012). These observations effectively point to a process that primarily operates in galaxy clusters, where star formation produces massive elliptical galaxies early on, followed by a quenching phase.
@@
Understanding how the large-scale environment led to a different star-formation process that resulted in such a stark differentiation between cluster galaxies and field galaxies is an open question. Many models have been put forward; on the one hand, galaxies could experience accelerated growth through increased merger rates in overdense environments (e.g., Kauffmann, 1996; Gottlöber et al., 2001; Fakhouri & Ma, 2009), while on the other hand gas could be removed from galaxies as they travel through the intracluster medium (ICM) due to ram pressure stripping (e.g., Gunn & Gott, 1972; Gavazzi et al., 2001; Boselli et al., 2019), or galaxies might be unable to accrete the new gas required to form more stars due to the hot ICM temperatures present in clusters (e.g., Larson et al., 1980; Balogh et al., 2000; Peng et al., 2015). To understand the contributions of these (and other) scenarios on galaxy evolution requires observations of cluster galaxies well before the cluster has fully virialized, which means finding examples of galaxy clusters early in their formation phase (i.e., protoclusters).
@@
Protoclusters are now common in the literature, having been found in large Lyman- surveys (e.g., Steidel et al., 2000; Chiang et al., 2015; Jiang et al., 2018), or around rare and extremely massive galaxies such as quasars (e.g., Dannerbauer et al., 2014; Noirot et al., 2018; Hennawi et al., 2015) and submillimetre (submm) galaxies (SMGs; e.g., Chapman et al. 2009; Casey et al. 2015; Oteo et al. 2018). However, such a diversity of selection techniques means that it is impossible to compare different protoclusters with one another and draw conclusions about overall populations.
@@
Nonetheless, cosmic microwave background (CMB) experiments are now making progress towards large, uniform samples of protoclusters. CMB surveys operate at submm and millimetre wavelengths over very large areas, inevitably detecting the brightest extragalactic foreground objects (relative to the CMB), many of which are now known to be protoclusters (e.g., Flores-Cacho et al., 2016; Miller et al., 2018; Koyama et al., 2021; Polletta et al., 2021). In particular, Planck has identified over 2000 extragalactic objects with spectral energy distributions (SEDs) peaking between 350 and 850 m (known as the Planck High- Catalogue, or PHz; see Planck Collaboration XXXIX 2016), meaning that the emission is most likely thermal and originating from copious amounts of dust produced by elevated star-formation rates (SFRs) in galaxies shortly before they quench.
@@
PLCK G073.457.5 (hereafter G073) is one such object. While G073 is not in the final PHz catalogue due to it being located just within the final Milky Way mask, it was part of a preliminary catalogue (with a less conservative mask) and was followed up with Herschel-SPIRE (Planck Collaboration XXVII, 2015) and SCUBA-2 (MacKenzie et al., 2017). Following these observations, ALMA was used to observe eight red SPIRE sources in Band 6 (1 mm), detecting 18 galaxies in the continuum (Kneissl et al., 2019). While 18 mm-bright galaxies is already 8–30 times higher than expected from a random field distribution, the distribution of photometric redshifts suggested that the field contained overdensities around and (although with large uncertainties), while two bright lines were also serendipitously detected at the edge of one of the sidebands and later identified as CO(5–4) at , coinciding with the first peak in the photometric redshift distribution.
@@
G073 therefore presents an opportunity to study overdense environments at a crucial epoch of cluster and galaxy formation. Around 2, galaxy clusters are beginning to virialize, and cluster galaxies often still have elevated SFRs compared to field galaxies (e.g., Hilton et al., 2010; Hayashi et al., 2011; Brodwin et al., 2013; Tran et al., 2015; Alberts et al., 2016; Nantais et al., 2017; Mei et al., 2023). For most of these systems, the star-formation activity is significant down to the dense core region (e.g., Hilton et al., 2010; Hayashi et al., 2010; Tran et al., 2010; Fassbender et al., 2011; Tadaki et al., 2012). Since star formation will often heavily obscure a galaxy in dust, these key populations can be missed by optical observations, but long-wavelength observations of molecular gas lines (primarily CO) are completely unaffected. Furthermore, since the molecular gas is the main fuel for star formation, such observations can be used to trace gas masses and provide insight into the evolutionary state of cluster and protocluster galaxies. A small number of (proto)cluster galaxies have been observed in CO around 2, with some results suggesting higher gas fractions () and shorter depletion timescales (SFR) compared to field galaxies (e.g., Noble et al., 2017; Hayashi et al., 2018), while others do not find such a difference (e.g., Aravena et al., 2012; Rudnick et al., 2017).
@@
The main issue to overcome is that our current observations are limited to only a few inhomogeneously-selected clusters and protoclusters. The PHz catalogue of over 2000 (uniformly-selected) protocluster candidates thus offers the opportunity to find statistically significant results with simple and understandable observational biases. To this end, in this paper we present a new ALMA survey of G073, searching for more galaxies and spectroscopically confirming the redshift of the structure. Ultimately, a full characterization of this object will help us understand how to carry out future studies of large samples of PHz objects.
@@
This paper is arranged as follows. In Sect. 2 we describe our new ALMA observations and all ancillary data, in Sect. 3 we outline our data analysis pipeline, and in Sect. 4 we discuss how we derive physical properties. Section 5 presents our results, which we interpret in Sect. 6. The paper concludes in Sect. 7. Throughout this work we adopt a Chabrier (2003) initial mass function (IMF) and a flat cold dark matter (CDM) model with cosmological parameters , , and km s-1 Mpc-1 (Planck Collaboration VI, 2020).
@@ 2.1 Target selection
Our ALMA Band 4 survey targeted the eight red SPIRE sources (in eight separate pointings) previously followed up by ALMA in Cycle 2 (Kneissl et al., 2019). These eight targets form a complete sample of SPIRE sources found within the Planck beam and with colours satisfying 250, 350 and 500 m flux density ratios and , which is expected to filter out low- interlopers. Since the ALMA primary beam is larger at 2 mm than at 1 mm, our observations covered all 18 galaxies detected by Kneissl et al. (2019). In addition to these eight SPIRE sources, we targeted the four highest signal-to-noise ratio (S/N) sources detected by SCUBA-2 at 850 m (again within the Planck beam; see MacKenzie et al. 2017), bringing the total number of ALMA pointings to 12. The distribution of these pointings is shown in Fig. 1.
@@ 2.2 Cycle 6 ALMA observations
In the LSB data, the redshift range covered by CO(3–2) was 1.49–1.56, while in the USB data Ci(1–0) covered a redshift range of 2.23–2.34. The lower sideband could also contain CO(4–3) between and , making a single line detection degenerate between the two expected structures; these situations are discussed on a case-by-case basis below. There are other potential contaminating lines such as CO(2–1) at and higher- lines at , which are strongly disfavoured given the photometric redshift distribution from Kneissl et al. (2019). There could additionally be Ci(1–0) in the lower sideband around , CO(3–2) in the upper sideband around , and CO(4–3) in the upper sideband around , but since multiple lines are detected in both the and structures (as shown below), we assume that there are only two redshift possibilities for all of the lines we detect. However for the SCUBA-2 targets we consider all possible redshifts because the original Planck selection would not have included objects bright at 850 m.
The Band 4 spectral set-up used four 1.78-GHz-wide continuum spectral windows (with channel widths of 15.6 MHz) around the central frequency of 143.1 GHz (2.1 mm). The spectral windows were divided into the two receiver sidebands, separated by 12 GHz (i.e., their central frequencies were 136.1, 138.1, 148.1, and 150.1 GHz). 45–48 antennas were available in the nominal C-5 array configuration with baseline lengths of 15–1400 m. This achieved synthesized beam sizes (using natural weighting) of around 0.65′′ 0.55′′(FWHM) in the lower side band and 0.58′′ 0.50′′in the upper sideband.
The observatory standard calibration was used. J00060623, a grid-monitoring source, was the bandpass and flux calibrator, with a flux density of 1.58 Jy at the central frequency. All pointings in the data sets shared the same phase calibrator, J23230317. The central continuum sensitivity in each of the 12 pointings was approximately 15–20 Jy beam-1. The reduction was based on the calibration provided by the ALMA Pipeline using standard CASA tasks (CASA Team et al., 2022), using natural weighting and pixel sizes of 0.077′′.
@@ 2.3 Cycle 2 ALMA observations
In this paper we make use of existing ALMA observations of G073 taken in Cycle 2. These observations were presented in detail in Kneissl et al. (2019); briefly, G073 was observed for 0.4 hours in Band 6 using the standard Band 6 tuning centred at about 1.3 mm. A total of eight pointings were carried out, targeting the same eight Herschel-SPIRE-detected sources described above (see Fig. 1). Based on these initial observations, a total of 18 individual galaxies were detected in the continuum, while line emission was detected in two of the 18 galaxies (IDs 3 and 8).
@@ 2.4 Ancillary data
G073 has been observed in the near-infrared (NIR; 1.25 m [] and 2.15 m []) with the Wide-field Infrared Camera (WIRCam) on the Canada France Hawaii Telescope and in the mid-infrared (MIR; 3.6 m and 4.5 m) with IRAC on the Spitzer Space Telescope. The observations and the photometric data are described in Kneissl et al. (2019). The NIR and MIR counterparts and fluxes for two ALMA sources, IDs 4 and 15, have been revised because they were wrongly assigned to a foreground galaxy located about 1.1′′and 1.6′′from the millimetre-emitting sources, respectively. Since ID 4 is much fainter than the nearby source, it is not possible to obtain reliable flux estimates for it. This is also true at sub-mm/mm wavelengths where the SED is consistent with a low redshift (i.e., ) source, suggesting that the nearby source dominates also at these wavelengths. The measured near and mid-IR flux densities , , , and are listed in Table 4.
@@ 3.1 CO and Ci line search for protocluster galaxies
To conduct our line search in both the lower and upper sideband of our ALMA observations, we first used the publicly-available tool LineSeeker (González-López et al., 2017, 2019). This takes as input a primary beam-uncorrected data cube, and convolves the cube along the spectral axis with Gaussians of varying width, searching for significant peaks. The noise per channel is estimated iteratively by computing the standard deviation of all the pixels in a given channel, then recomputing the standard deviation of all the pixels whose absolute values are lower than 5 times the initial standard deviation. After convolution with a Gaussian of a given size, pixels whose spectra show peak S/N values above a chosen threshold are then returned to the user, and the procedure is repeated to search for pixels containing negative S/N peaks.
@@
From the catalogues of positive and negative peaks returned by LineSeeker, we found that the most significant negative peak across all of the data cubes had an S/N of 6.1, thus we used this as our cutoff for identifying real positive peaks. There were a total of eight unique positive peaks in the LSB data with S/N values greater than 6.1; six of these peaks are spatially coincident with Band 6 continuum-detected galaxies from Kneissl et al. (2019), one peak is located in a field previously observed by ALMA but at a position with no continuum counterpart (ID 18), and one peak is in the S2-3 pointing (ID 24; see Fig. 1). Similarly, we found three unique positive peaks in the USB data above the S/N threshold, with two peaks coincident with a Band 6 continuum-detected galaxy, and one peak with no previously known counterpart (ID 19).
For completeness, we also placed 1.5′′ apertures around the previously-detected galaxies in G073 from Kneissl et al. (2019) with no line detections from LineSeeker. Since for these galaxies we have prior knowledge of their positions, we systematically search their spectra for fainter peaks down to 5 (without any Gaussian convolution), but this did not return any additional lines.

### v2

@@
Planck’s High-Frequency Instrument observed the whole sky between 350 m and 3 mm, discovering thousands of unresolved peaks in the cosmic infrared background. The nature of these peaks is still poorly understood – while some are strong gravitational lenses, the majority are overdensities of star-forming galaxies but with almost no redshift constraints. PLCK G073.457.5 (G073) is one of these Planck-selected peaks. ALMA observations of G073 suggest the presence of two structures (one around redshift 1.5 and one around redshift 2) aligned along the line of sight, but this analysis lacked robust spectroscopic confirmation. Characterizing the full redshift distribution of the galaxies within G073 is needed in order to better understand this representative example of Planck-selected objects, and connect them to the emergence of galaxy clusters. We used ALMA Band 4 spectral scans to search for CO(3–2), CO(4–3), and Ci(1–0) line emission, targeting eight red Herschel-SPIRE sources in the field, as well as four bright SCUBA-2 sources. We find 15 emission lines in 13 galaxies, and using existing photometry information, we determined the spectroscopic redshift of all 13 galaxies. Eleven of these galaxies are SPIRE-selected and lie in two structures at and (both with a standard deviation in redshift of 0.02), while the two SCUBA-2-selected galaxies are at . Using multi-wavelength photometry, we constrained stellar masses and star formation rates, and using the CO and Ci emission lines we constrained gas masses. Our protocluster galaxies exhibit typical depletion timescales (/SFR) for field galaxies at the same redshifts but higher gas-to-stellar mass ratios, potentially driven by emission line selection effects. The two structures confirmed in our survey are reproduced in cosmological simulations of star-forming halos at high redshifts; the simulated halos have a 60–70% probability of collapsing into galaxy clusters, implying that the two structures in G073 are genuinely protoclusters.
@@
The Universe on the largest scales forms a ‘cosmic web’, with galaxy clusters at the intersections of the filaments (Bond et al. 1996). These are the largest gravitationally bound objects in the Universe, and we believe that they play an important role in the evolution of the galaxies embedded within them because in the local Universe, cluster galaxies are more likely to be massive, red ellipticals with old stellar populations (e.g., Ellis et al. 1997; Andreon 2003; Muzzin et al. 2012). These observations effectively point to a process that primarily operates in galaxy clusters, where star formation produces massive elliptical galaxies early on, followed by a quenching phase.
@@
Understanding how the large-scale environment led to a different star formation process that resulted in such a stark differentiation between cluster galaxies and field galaxies is an open question. Many models have been put forward; on the one hand, galaxies could experience accelerated growth through increased merger rates in overdense environments (e.g., Kauffmann 1996; Gottlöber et al. 2001; Fakhouri & Ma 2009), while on the other hand gas could be removed from galaxies as they travel through the intracluster medium due to ram pressure stripping (e.g., Gunn & Gott 1972; Gavazzi et al. 2001; Boselli et al. 2019), or galaxies might be unable to accrete the new gas required to form more stars due to the hot intracluster medium temperatures present in clusters (e.g., Larson et al. 1980; Balogh et al. 2000; Peng et al. 2015). To understand the contributions of these (and other) scenarios to galaxy evolution requires observations of cluster galaxies well before the cluster has fully virialized, which means finding examples of galaxy clusters early in their formation phase (i.e. protoclusters).
@@
Protoclusters are now common in the literature, having been found in large Lyman- surveys (e.g., Steidel et al. 2000; Chiang et al. 2015; Jiang et al. 2018), or around rare and extremely massive galaxies such as quasars (e.g., Dannerbauer et al. 2014; Noirot et al. 2018; Hennawi et al. 2015) and submillimetre (submm) galaxies (SMGs; e.g., Chapman et al. 2009; Casey et al. 2015; Oteo et al. 2018). However, such a diversity of selection techniques means that it is impossible to compare different protoclusters with one another and draw conclusions about overall populations.
@@
Nonetheless, thanks to cosmic microwave background (CMB) experiments large, uniform samples of protoclusters are being compiled. CMB surveys operate at submm and millimetre wavelengths over very large areas, inevitably detecting the brightest extragalactic foreground objects (relative to the CMB), many of which are now known to be protoclusters (e.g., Flores-Cacho et al. 2016; Miller et al. 2018; Koyama et al. 2021; Polletta et al. 2021). In particular, Planck has identified over 2000 extragalactic objects with spectral energy distributions (SEDs) peaking between 350 and 850 m, meaning that the emission is most likely thermal and originating from copious amounts of dust produced by elevated star-formation rates (SFRs) in galaxies shortly before they quench; these objects are listed in the Planck High- (PHz) catalogue (Planck Collaboration XXXIX 2016).
@@
PLCK G073.457.5 (hereafter G073) is one such object. While G073 is not in the final PHz catalogue due to being located just within the final Milky Way mask, it was part of a preliminary catalogue (with a less conservative mask) and was followed up with the Herschel Spectral and Photometric Imaging REceiver (SPIRE; Planck Collaboration XXVII 2015) and the Submillimetre Common-User Bolometer Array 2 (SCUBA-2; MacKenzie et al. 2017). Following these observations, the Atacama Large Millimeter/submillimeter Array (ALMA) was used to observe eight red SPIRE sources in Band 6 (1 mm), detecting 18 galaxies in the continuum (Kneissl et al. 2019). While 18 millimetre-bright galaxies is already 8–30 times higher than expected from a random field distribution, the distribution of photometric redshifts suggests that the field contained overdensities around and (although with large uncertainties), while two bright lines were also serendipitously detected at the edge of one of the sidebands and later identified as CO(5–4) at , coinciding with the first peak in the photometric redshift distribution.
@@
G073 therefore presents an opportunity to study overdense environments at a crucial epoch of cluster and galaxy formation. Around 2, galaxy clusters are beginning to virialize, and cluster galaxies often still have elevated SFRs compared to field galaxies (e.g., Hilton et al. 2010; Hayashi et al. 2011; Brodwin et al. 2013; Tran et al. 2015; Alberts et al. 2016; Nantais et al. 2017; Mei et al. 2023). For most of these systems, the star formation activity is significant down to the dense core region (e.g., Hilton et al. 2010; Hayashi et al. 2010; Tran et al. 2010; Fassbender et al. 2011; Tadaki et al. 2012). Since star formation will often heavily obscure a galaxy in dust, these key populations can be missed by optical observations, but long-wavelength observations of molecular gas lines (primarily CO) are completely unaffected. Furthermore, since the molecular gas is the main fuel for star formation, such observations can be used to trace gas masses and provide insight into the evolutionary state of cluster and protocluster galaxies. A small number of (proto)cluster galaxies have been observed in CO around 2, with some results suggesting higher gas fractions () and shorter depletion timescales (SFR) compared to field galaxies (e.g., Noble et al. 2017; Hayashi et al. 2018), while others do not find such a difference (e.g., Aravena et al. 2012; Rudnick et al. 2017).
@@
The main issue to overcome is that our current observations are limited to only a few inhomogeneously selected clusters and protoclusters. The PHz catalogue of over 2000 (uniformly selected) protocluster candidates thus offers the opportunity to find statistically significant results with simple and understandable observational biases. To this end, we present a new ALMA survey of G073, searching for more galaxies and spectroscopically confirming the redshift of the structure. Ultimately, a full characterization of this object will help us understand how to carry out future studies of large samples of PHz objects.
@@
This paper is arranged as follows. In Sect. 2 we describe our new ALMA observations and all ancillary data, in Sect. 3 we outline our data analysis pipeline, and in Sect. 4 we discuss how we derived physical properties. Section 5 presents our results, which we interpret in Sect. 6. The paper concludes in Sect. 7. Throughout this work we adopt a Chabrier (2003) initial mass function (IMF) and a flat cold dark matter (CDM) model with cosmological parameters , , and km s-1 Mpc-1 (Planck Collaboration VI 2020).
@@ 2.1 Target selection
Our ALMA Band 4 survey targeted the eight red SPIRE sources (in eight separate pointings) previously followed up by ALMA in Cycle 2 (Kneissl et al. 2019). These eight targets form a complete sample of SPIRE sources found within the Planck beam and with colours satisfying 250, 350 and 500 m flux density ratios and , which is expected to filter out low- interlopers. Since the ALMA primary beam is larger at 2 mm than at 1 mm, our observations covered all 18 galaxies detected by Kneissl et al. (2019). In addition to these eight SPIRE sources, we targeted the four highest signal-to-noise ratio (S/N) sources detected by SCUBA-2 at 850 m (again within the Planck beam; see MacKenzie et al. 2017), bringing the total number of ALMA pointings to 12. The distribution of these pointings is shown in Fig. 1.
@@ 2.2 Cycle 6 ALMA observations
In the LSB data, the redshift range covered by CO(3–2) was 1.49–1.56, while in the USB data Ci(1–0) covered a redshift range of 2.23–2.34. The LSB could also contain CO(4–3) between and , making a single line detection degenerate between the two expected structures; these situations are discussed on a case-by-case basis below. There are other potential contaminating lines such as CO(2–1) at and higher- lines at , which are strongly disfavoured given the photometric redshift distribution from Kneissl et al. (2019). There could additionally be Ci(1–0) in the LSB around , CO(3–2) in the USB around , and CO(4–3) in the USB around , but since multiple lines are detected in both the and structures (as shown below), we assumed that there are only two redshift possibilities for all of the lines we detected. However, for the SCUBA-2 targets we considered all possible redshifts because the original Planck selection would not have included objects bright at 850 m.
The Band 4 spectral set-up used four 1.78-GHz-wide continuum spectral windows (with channel widths of 15.6 MHz) around the central frequency of 143.1 GHz (2.1 mm). The spectral windows were divided into the two receiver sidebands, separated by 12 GHz (i.e., their central frequencies were 136.1, 138.1, 148.1, and 150.1 GHz). 45–48 antennas were available in the nominal C-5 array configuration with baseline lengths of 15–1400 m. This achieved synthesized beam sizes (using natural weighting) of around 0.65′′ 0.55′′full width at half maximum (FWHM) in the lower side band and 0.58′′ 0.50′′in the USB.
The observatory standard calibration was used. J00060623, a grid-monitoring source, was the bandpass and flux calibrator, with a flux density of 1.58 Jy at the central frequency. All pointings in the datasets shared the same phase calibrator, J23230317. The central continuum sensitivity in each of the 12 pointings was approximately 15–20 Jy beam-1. The reduction was based on the calibration provided by the ALMA Pipeline using standard CASA tasks (CASA Team et al. 2022), using natural weighting and pixel sizes of 0.077′′.
@@ 2.3 Cycle 2 ALMA observations
We made use of existing ALMA observations of G073 taken in Cycle 2. These observations are presented in detail in Kneissl et al. (2019); briefly, G073 was observed for 0.4 hours in Band 6 using the standard Band 6 tuning centred at about 1.3 mm. A total of eight pointings were carried out, targeting the same eight Herschel-SPIRE-detected sources described above (see Fig. 1). Based on these initial observations, a total of 18 individual galaxies were detected in the continuum, while line emission was detected in two of the 18 galaxies (IDs 3 and 8).
@@ 2.4 Ancillary data
G073 has been observed in the near-infrared (NIR; 1.25 m [] and 2.15 m []) with the Wide-field Infrared Camera (WIRCam) on the Canada France Hawaii Telescope and in the mid-infrared (MIR; 3.6 m and 4.5 m) with the Infrared Array Camera (IRAC) on the Spitzer Space Telescope. The observations and the photometric data have already been presented and described in Kneissl et al. (2019). The NIR and MIR counterparts and fluxes for two ALMA sources, IDs 4 and 15, have been revised because they were wrongly assigned to a foreground galaxy located about 1.1′′and 1.6′′from the millimetre-emitting sources, respectively. Since ID 4 is much fainter than the nearby source, it is not possible to obtain reliable flux estimates for it. This is also true at submm/millimetre wavelengths where the SED is consistent with a low redshift (i.e. ) source, suggesting that the nearby source dominates also at these wavelengths. The measured NIR and MIR flux densities , , , and are listed in Table 4.
Since multiple ALMA detections are often found within a single SPIRE source, we assigned de-blended submm flux densities as done in Kneissl et al. (2019). The redshifts used in the de-blending procedure are all consistent with the revised redshifts (see Sect. 3), with the exception of ID 9 for which the revised redshift is smaller (i.e. compared to the previous value, ). Since this source is detected at 850 m and 1.2 mm, and the de-blended flux density estimates are derived only for 350 m and 500 m (this source was not detected at 250 m), we assumed that its far-infrared (FIR) SED and estimated FIR luminosity are not significantly affected by the SPIRE flux densities derived from the de-blending procedure using a slightly higher redshift.
@@ 3.1 CO and Ci line search for protocluster galaxies
To conduct our line search in both the LSB and USB of our ALMA observations, we first used the publicly available tool LineSeeker (González-López et al. 2017, 2019). This takes as input a primary beam-uncorrected data cube, and convolves the cube along the spectral axis with Gaussians of varying width, searching for significant peaks. The noise per channel is estimated iteratively by computing the standard deviation of all the pixels in a given channel, then recomputing the standard deviation of all the pixels whose absolute values are lower than 5 times the initial standard deviation. After convolution with a Gaussian of a given size, pixels whose spectra show peak S/N values above a chosen threshold are then returned to the user, and the procedure is repeated to search for pixels containing negative S/N peaks.
@@
From the catalogues of positive and negative peaks returned by LineSeeker, we found that the most significant negative peak across all of the data cubes had an S/N of 6.1; thus, we used this as our cutoff for identifying real positive peaks. There were a total of eight unique positive peaks in the LSB data with S/N values greater than 6.1; six of these peaks are spatially coincident with Band 6 continuum-detected galaxies from Kneissl et al. (2019), one peak is located in a field previously observed by ALMA but at a position with no continuum counterpart (ID 18), and one peak is in the S2-3 pointing (ID 24; see Fig. 1). Similarly, we found three unique positive peaks in the USB data above the S/N threshold, with two peaks coincident with a Band 6 continuum-detected galaxy, and one peak with no previously known counterpart (ID 19).
For completeness, we also placed 1.5′′ apertures around the previously detected galaxies in G073 from Kneissl et al. (2019) with no line detections from LineSeeker – this aperture size was chosen to be slightly smaller than the average aperture size constructed for the detected sources, as described in Sect. 3.3. Since for these galaxies we have prior knowledge of their positions, we systematically searched their spectra for fainter peaks down to 5 (without any Gaussian convolution), but this did not return any additional lines.